The mean potassium content of a popular sports drink is listed as 149 mg in a 32-oz bottle. Analysis of 28 bottles indicates a sample mean of 148 mg.

(a) State the hypotheses for a two-tailed test of the claimed potassium content.
a. H0: μ = 149 mg vs. H1: μ ≠ 149 mg
b. H0: μ ≤ 149 mg vs. H1: μ > 149 mg
c. H0: μ ≥ 149 mg vs. H1: μ < 149 mg
(b)

Assuming a known standard deviation of 2.2 mg, calculate the z test statistic to test the manufacturer’s claim. (Round your answer to 2 decimal places. A negative value should be indicated by a minus sign.)

(c)

At the 10 percent level of significance (α = 0.1) does the sample contradict the manufacturer’s claim?

  

Decision Rule: Reject H0     (Click to select)   

a) if z > + 1.645 or if z < -1.645   

b) if z < + 1.645 or if z < -1.645  

c) if z < + 1.645 or if z > -1.645

  

The sample   (Click to select) the manufacturer's claim.

a) contradicts   

b) does not contradict

(d)

Find the p-value. (Round intermediate calculations to 2 decimal places. Round your answer to 4 decimal places.)

  p-value   
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

b)

= (148 - 149)/2.2/sqrt(28) = -2.4052

C)

For 90% confidence interval z value is

Decision rule =

Reject H0 if z < -1.645 or z > 1.645

since calculated z lesser than -.1645, we reject the hypothesis

d)

P-value = 2 * P( Z < z)

= 2 * P( Z < -2.4052)

=2*0.008082

= 0.016164